# Data Distribution-based Curriculum Learning

### v1

@@ VI Conclusion
This paper proposed a curriculum learning approach known as Data Distribution-based Curriculum Learning (DDCL). It used the data distribution of a dataset to build a curriculum based on the order of samples. Two types of scoring methods known as DDCL (Density) and DDCL (Point) were used to score samples thus determining their training order. DDCL (Density) used the sample density to assign scores while DDCL (Point) utilised the Euclidean distance for scoring.
Experiments were conducted on multiple datasets using a neural network, SVM and random forest classifier to evaluate DDCL. Results showed that while performance varied across the datasets depending on the classifier used, the application of DDCL resulted in significant increases in the average accuracy for most datasets compared to not using any curriculum. Furthermore, analysis of the error losses for five training epochs using batch gradient descent reveals that convergence is faster when using DDCL with either of the scoring methods over the no curriculum method.
@@
Future work will explore the creation of additional scoring methods to determine their impact on the training performance and investigate the application of ensemble learning using each scoring method. In addition, self-paced learning concepts will be incorporated into DDCL in order to dynamically determine the curriculum based on feedback from the learner. The proposed approach can also be adapted to image and text processing applications.

### v2

@@ VI Conclusion
This paper proposed a curriculum learning approach known as Data Distribution-based Curriculum Learning (DDCL). It used the data distribution of a dataset to build a curriculum based on the order of samples. Two types of scoring methods known as DDCL-Density and DDCL-Point were used to score samples thus determining their training order. DDCL-Density used the sample density to assign scores whereas DDCL-Point utilized the Euclidean distance for scoring.
Experiments were conducted on multiple medical datasets using a neural network, SVM and random forest classifier to evaluate DDCL. Results showed that even though performance varied across the datasets between the classifiers used, the application of DDCL resulted in accuracy increases ranging from 2% to 10% for all datasets compared to other state-of-the-art methods. Furthermore, analysis of the error losses for five training epochs using batch gradient descent reveals that convergence is faster when using DDCL with either of the scoring methods over the no curriculum method.
@@
Future work will explore the creation of additional scoring methods to determine their impact on the training performance and investigate the application of ensemble learning using each scoring method. By introducing these new scoring methods, we aim to assess their effectiveness in comparison to other curriculum approaches such as SPL and Reverse Curriculum Learning (RCL). In addition, self-paced learning concepts will be incorporated into DDCL in order to dynamically determine the curriculum based on feedback from the learner. This integration will allow the model to adjust its learning pace according to its current performance, potentially leading to a more robust and adaptive training process. Furthermore, the algorithms of the proposed DDCL’s scoring methods suggests versatility which can be adapted beyond its initial application to structured datasets. Specifically, the proposed approach can also be adapted to image and text processing applications where the inherent complexity and variability of data types present unique challenges.
Nomenclature
Symbol Description
$S$ Grouped data samples for all classes
$s$ Data for a specific class
${O_{s}}$ Centroids for a specific class
$E_{s}$ Euclidean distances from centroid (dimensionless)
$D_{s}$ Data distribution for a class
$Q^{s}$ Quantiles for the whole dataset
$q$ A single quantile in a dataset
$R_{s}$ Scored training data
$T$ Curriculum prepared training data
$C_{s}$ Densities (count) for all quantiles
$\hat{E_{s}}$ Normalized Euclidean distances from centroid